OpenStudy (anonymous):

Which of the following is the slope between the two points (-3, 1) and (3, -6)?

OpenStudy (anonymous):

right 6 units (from -3 to 3) down seven units (from 1 to -6) slope is \(-\frac{7}{6}\)

OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? Choose one answer. A. 2x + 4y = 8 B. 4y = -2x + 8 C. y = -1/2x + 2 D. x = -2y + 4